Parth (parthkohli):

I'd now love to congratulate the super-five who just joined us :) @Diyadiya @apoorvk @Loujoelou @saifoo.khan @Akshay_Budhkar Congrats!

Parth (parthkohli):

\(\Huge \mathtt{\text{Welcome to our team!}}\)

Parth (parthkohli):

$HAKA~~

OpenStudy (lgbasallote):

hmm..feels like someone is missing

OpenStudy (diyadiya):

\[\Large \mathbf{\mathcal {THANK ~ YOU ~~ =)}}\]

OpenStudy (lgbasallote):

tomas a.

mathslover (mathslover):

Congrats from my side too: \[\Huge{\mathbb{C}\textbf{O}\textit{N}\mathbb{G}\textbf{R}\mathbb{A}\textbf{T}\mathbb{S}}\] \begin{array}l\color{red}{\normalsize\text{C}}\color{orange}{\normalsize\text{o}}\color{#9c9a2e}{\normalsize\text{n}}\color{green}{\normalsize\text{g}}\color{blue}{\normalsize\text{r}}\color{purple}{\normalsize\text{a}}\color{purple}{\normalsize\text{t}}\color{red}{\normalsize\text{s}}\color{orange}{\normalsize\text{ }}\color{#9c9a2e}{\normalsize\text{t}}\color{green}{\normalsize\text{o}}\color{blue}{\normalsize\text{ }}\color{purple}{\normalsize\text{a}}\color{purple}{\normalsize\text{l}}\color{red}{\normalsize\text{l}}\color{orange}{\normalsize\text{ }}\color{#9c9a2e}{\normalsize\text{t}}\color{green}{\normalsize\text{h}}\color{blue}{\normalsize\text{e}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{n}}\color{red}{\normalsize\text{e}}\color{orange}{\normalsize\text{w}}\color{#9c9a2e}{\normalsize\text{ }}\color{green}{\normalsize\text{a}}\color{blue}{\normalsize\text{m}}\color{purple}{\normalsize\text{b}}\color{purple}{\normalsize\text{a}}\color{red}{\normalsize\text{s}}\color{orange}{\normalsize\text{s}}\color{#9c9a2e}{\normalsize\text{a}}\color{green}{\normalsize\text{d}}\color{blue}{\normalsize\text{o}}\color{purple}{\normalsize\text{r}}\color{purple}{\normalsize\text{s}}\color{red}{\normalsize\text{}}\end{array}

OpenStudy (lgbasallote):

dgo

Parth (parthkohli):

dgo?

Parth (parthkohli):

He's no more an amby ash2326 sasogeek 2bornot2b Zed GOODMAN RahulZ Tomas.A rebeccaskell94 karatechopper ParthKohli lgbasallote mathslover Loujoelou Diyadiya apoorvk saifoo.khan Akshay_Budhkar

Parth (parthkohli):

I added Zed anyway.

OpenStudy (lgbasallote):

so what was the list for again?

mathslover (mathslover):

:D supa 16 now ?

Parth (parthkohli):

lol

OpenStudy (apoorvk):

\[\huge \mathbb{\color{green}{\text{THANK YOU }}}\mathbb{\color{#D4A017}{\text{SIRE!! }}}\color{#8417DD}{ツ}\]

OpenStudy (akshay_budhkar):

Thank you!!!! :D Would love to work along you people :)

OpenStudy (anonymous):

Hey everyone! Wow, so many more ambys! OS will never fall. Hahaha. Congrats everyone!

OpenStudy (carniel):

I think the list isnt complete e.e Colm should add me.

OpenStudy (unklerhaukus):

you peoples are great peoples

OpenStudy (karatechopper):

WOOT WOOT!

OpenStudy (anonymous):

Congrats !!!!!!!!!!! to all the new Ambassadors ! wish you to be OS favorites . And to all Old and First Ambassadors , ( yay ) we have cool friends now !